Session: Feb 2016 (session 1)
Wavelength: 6 cm
To give feedback on this experiment, click
(Mariotti at Mon Feb 22 15:13:41 2016) Success:
some RFI on band
(PENG LI at Mon Feb 22 15:53:48 2016) Minor failures:
UT 01: 00-09: 45, VEX files generated PRC file to decimate the
parameters for errors, modify the parameters of the experiment returned
to normal after the decimate.
(Bezborodov A. at Mon Feb 22 19:32:59 2016) Success:
(Plotnikova, Zhukova, Lysakova at Mon Feb 22 19:33:14 2016) Success:
(Shumilova at Mon Feb 22 19:34:07 2016) Minor failures: RECO
Scans No0001-N0010 failed due to Mark5B+ problems.
(Daniel Valero at Tue Feb 23 08:29:42 2016) Minor failures:
wrong sync until 4:00 UTC
(Jun Yang at Tue Feb 23 13:53:19 2016) Success:
Stowed the telescope due to high winds during 3:40 - 6:02 UT.
(Geert Kuper at Tue Feb 23 22:35:25 2016) Success: FS
Field system was stopped by accident. Schedule stopped at 10:30 UT and was restarted at 10:38. Recording resumed at 10:48
(Uwe Bach at Wed Feb 24 08:53:49 2016) Minor failures: WIND
First half of the time was lost due to strong wind. Observations at Ef started at 8 UT.
(Geert Kuper at Wed Feb 24 10:06:25 2016) Success:
Flagr software sometimes slow. The RT is mostly earlier on position as flagr indicates.
(P.Cassaro at Thu Feb 25 13:22:52 2016) Minor failures: TSYS
The experiment has been successfully recorded, but the tsys are too high during the whole experiment, due to receiver cooling problems.
(Pawel Wolak at Fri Feb 26 14:57:51 2016) Success:
(Bo Xia at Sat Feb 27 04:45:20 2016) Minor failures:
Because of the trouble of the power supply before the session
started, we found the temperature of the receiver was returning.
At UT 7:17, we modified the decimation ratio of the prc file to the
correct values. The formatter clock has changed for 2us. So we reloaded
the system and recooling. All system was nomal at UT 9:00.
(Bo Xia at Sat Feb 27 04:47:03 2016) Failure:
Didn't observe dut to the maintenance.